# Automated Test Transfer Across Android Apps Using Large Language Models

### v1

@@ 4.1. Experimental Setup
5. Threats to Validity and Limitations
@@
Similar to the existing test transfer techniques, the performance of LLMigrate relies on the accessibility of the target app. As the proposed technique utilizes the XML layout hierarchy of each app state to generate appropriate events, its performance can be compromised if the target app’s XML hierarchy layout lacks meaningful content. Specifically, the absence of descriptive attribute values such as resource-id or content-desc for identifying elements’ functionality can adversely impact LLMigrate’s effectiveness. Recent research work (Chen et al., 2020) has demonstrated promising results for generating descriptive attributes for elements from their visual content. These approaches can be utilized in future studies to improve the accessibility of the target before starting the test transfer.
Certain app elements containing important information, such as toast messages, may only be visible for a short duration due to their transient nature. This poses two significant challenges for the tool. First, an event targeting such a dynamic element may be created, but when attempting to execute and verify the event, the element may have disappeared from the page as the tool lacks the capability to freeze the app in a fixed state. Second, the technique may fail to capture the content relevant to the transient element in time, which leads to a failure in creating the proper event.
Some events are irreversible, and cannot be repaired even with a complete restart. For example, in the case of transferring a registration test, if LLMigrate encounters an exception in the transfer after the step of entering email and password, the tool may not be able to backtrack to re-register using the same email and password, as it may encounter server errors. In general, maintaining the app state outside the app, such as on a server, can lead to difficulties in test transfer.
@@ 6. Related Works
The most relevant group of works targets transferring tests from one Android app to another, similar to LLMigrate. Behrang et al. (Behrang and Orso, 2018, 2019) and Lin et al. (Lin et al., 2019b) proposed ATM and CraftDroid, which rely on app analysis and NLP techniques to transfer tests across different Android apps within the same domain. AppFlow (Hu et al., 2018) is a machine learning-based approach that utilized screen and widget classification to generate UI tests for an app using a library of existing tests. Liu et al. (Liu et al., 2022, 2024d) proposed adaptive semantic matching strategies for test transfer. Yu et al. (Zhang et al., 2024b) have recently proposed TEMdroid, a semantic matching-based approach for test transfer that leverages dynamic analysis and Siamese networks. Zhao et al. (Zhao et al., 2020) proposed FrUITeR, a framework for automatically evaluating the previous test transfer approaches. Mariani et al. (Mariani et al., 2021) presented a study on techniques for semantic matching of GUI events used by existing test reuse approaches. Khalili et al. proposed SemFinder (Khalili et al., 2024), an approach that assesses different configurations for mapping UI events across apps based on their textual information but does not focus on the test transfer problem as a whole. Mishra et al. (Mishra et al., 2023) extended SemFinder by incorporating visual information into the event mapping technique.
@@
Another group of recent works targets a slightly different problem, focusing on many-to-one UI test transfer across Android apps. MigratePro (Zhang et al., 2024c) is a technique that aims to improve UI test transfer by generating a new test from multiple tests that have already been migrated to the target app from various source apps. MigratePro is not a transfer technique itself, and it improves tests transferred by an existing transfer technique. Future research can explore using tests transferred by LLMigrate as input for MigratePro to assess potential improvements. Another recent work, MACdroid (Zhang et al., 2024a), uses LLMs to create tests for a new target app using an abstract test logic created from multiple source tests targeting the same functionality on different apps. Note that compared to one-to-one transfer, many-to-one transfer is a less challenging problem due to the availability of more data, such as multiple different flows in various apps that can be more similar to the intended flow in the target app. However, in practice, there often are not multiple compatible tests for the same functionality available that can be used for transfer, which limits the practicality of these techniques. Furthermore, unlike LLMigrate, which employs multimodal LLMs and utilizes both visual and textual information for UI understanding task, MACdroid relies solely on textual data, which can lead to certain limitations, as discussed in Section 3. A more detailed comparison of the two approaches is not possible due to the unavailability of the implementation and artifacts relevant to this technique at the time of publication.
@@
Finally, several recent publications (Kang et al., 2023b, a; Feng and Chen, 2024; Huang et al., 2024; Liu et al., 2024a, b, 2023; Yu et al., 2023; Yoon et al., 2023; Wen et al., 2023; Ju et al., 2024) have explored the application of LLMs to advance mobile testing. However, none of these techniques address the automated transfer of existing usage-based tests across apps with similar functionality.
@@ 7. Conclusion and Future Work
This paper presents LLMigrate, a technique that relies on multimodal LLMs for transferring usage-based UI tests across Android apps. In our extensive evaluation covering five app categories, LLMigrate was able to successfully transfer 97% of tests and reduce more than 90% of the total manual work required for writing UI tests.
Potential future areas of work can target expanding our technique, particularly for development across various platforms such as Web and iOS, which promises to yield significant time savings in E2E test development and maintenance. Another possible research direction is the study of how integrating the approaches (Chen et al., 2020) that enhance app accessibility can improve the applicability of techniques such as LLMigrate on non-accessible apps.
@@ 8. Data Availability
LLMigrate’s implementation and all of our research artifacts are available publicly (rep, [n. d.]).

### v2

@@ 4.1. Experimental Setup
4.6. RQ5. Evaluating LLMigrate on Data Unseen by LLMs
Since LLMigrate leverages existing off-the-shelf LLMs like GPT-4o to transfer tests, it is important to evaluate its effectiveness on apps and test cases that the LLM has not previously encountered. This helps ensure that the promising results discussed earlier are not simply due to the LLM’s prior familiarity with the subjects, and it also validates LLMigrate ’s applicability to future, unseen apps. To do this, we took two key steps: selecting apps that were released after GPT-4o’s knowledge cutoff date of October 2023 (OpenAI, [n. d.]), and manually writing new test cases for these apps rather than relying on publicly available ones. We introduced five new app categories and, within each, selected three recently released apps. For each category, we manually designed a representative usage scenario and created corresponding tests, resulting in 30 total test transfers across these new apps.
Table 5 outlines the app categories, selected apps, and usage scenarios. Table 6 presents the precision, recall, success rate, reduction, and transfer time metrics for these unseen test transfers. On average, LLMigrate achieved 97% and 89% precision and 97% and 92% recall for GUI and oracle events, respectively. Additionally, it reached a 93% average successful transfer rate and 88% reduction score, consistent with the findings in earlier research questions. These outcomes confirm that LLMigrate performs reliably on data that was not part of the LLM’s training set, supporting the robustness of the approach.
Table 5. Evaluation subjects used to verify LLMigrate’s applicability on data previously unseen to GPT-4o.
App Category Subject Apps Release Functionality # Total # Oracle
Date Events Events
c6-EMI Calculator a1-EMI Calculator & Financial (emi, 2025b) 9 Aug 2024 t1-Add 3 entries and calculate the EMI 5.4 1
a2-Cash Loan EMI Calculator (emi, 2025a) 20 Jan 2025
a3-EMI Calculator : Loan Planner (emi, 2025c) 1 Feb 2025
c7-AI Chatbots a1-Google Gemini (cha, 2025c) 4 Jun 2024 t1-Start a conversation asking a Yes/No question 3 2
a2-Deep Search - AI Chatbot (cha, 2025b) 30 Jan 2025 and verify the answer
a3-Chatbot - AI Smart Assistant (cha, 2025a) 16 Feb 2024
c8-Movies a1-Ava Assistant - Movies & Shows (mov, 2025b) 29 Jan 2025 t1-Search for a movie and share the details 4.5 1
a2-200TV - Live TV Movies App (mov, 2025a) 16 Jan 2025
a3-ClipFix: Movie Shazam (mov, 2025c) 29 Aug 2024
c9-Note a1-Daily Notes - Easy Notebook (not, 2025a) 1 Feb 2025 t1-Add a note and search for it by its title 6.5 2
a2-Notes - QuickNotes (not, 2025b) 14 Jan 2024
a3-Personal notes and tasks (not, 2025c) 5 Jan 2025
c10-Messenger a1-Messages for SMS - DUAL SIM (sms, 2025b) 26 Dec 2024 t1-Search for a phone number and send a message 5 2
a2-Messages: Text SMS (sms, 2025c) 29 Jan 2025 to the found recipient
a3-Color SMS: Message & Messenger (sms, 2025a) 26 Oct 2023
Table 6. Analysis of precision, recall, success, reduction, and transfer time metrics achieved by LLMigrate, on unseen apps.
Category Test Precision Recall Successful Reduction Transfer Time
GUI Event Oracle Event GUI Event Oracle Event Transfer Rate
c6-t1 100 100 100 100 100 100 294.5
c7-t1 100 66.7 100 72.7 100 60 265.44
c8-t1 100 30 100 100 100 100 273.69
c9-t1 91.7 91.7 91.7 91.7 66.7 83.3 333.77
c10-t1 100 100 100 100 100 100 424.63
Average 97.8 89.6 97.8 92.8 93.3 88.6 318.41
5. Discussion
Our evaluation demonstrates that LLMigrate achieves high accuracy in transferring tests across Android apps, outperforming existing solutions. However, we observed certain failure cases that highlight limitations of the current approach. These include incorrect action selection by the LLM, incomplete flow transfer where critical steps are omitted, and the insufficient oracles generation, particularly for transitions. While these issues do not always prevent test execution, they can compromise the correctness or completeness of the transferred tests. Readers can find specific examples of these cases on the project repository (rep, 2025).
Additionally, LLMigrate shares some common limitations with prior test transfer techniques. Its performance is influenced by the accessibility and quality of UI metadata in the target app, particularly when meaningful attributes like resource-id or content-desc are missing. The tool also struggles with transient UI elements such as toast messages, which may disappear before they can be processed. Furthermore, certain events, like user registration, can be irreversible, making it difficult to recover from partial failures. These challenges suggest directions for future enhancements, such as incorporating visual analysis (Chen et al., 2020) or improving app state management.
6. Threats to Validity
@@
Another threat to the validity of our evaluation comes from our inability to run the previous test transfer tools. This is due to reasons such as the unavailability of parts of the source code (Zhang et al., 2024a) and technical problems that are mainly due to changes and the lack of maintenance of existing dependencies (Lin et al., 2019b; Liu et al., 2024b). As detailed in Section 4, to navigate this issue and offer a comparative analysis, we utilized the publicly available CraftDroid dataset (Lin et al., 2019b) previously used by existing tools. A related threat arises from a small number of apps from the CraftDroid dataset (Lin et al., 2019b) that are deprecated. To mitigate this threat and ensure a fair comparison, for CraftDroid (Lin et al., 2019b) and TREADROID (Liu et al., 2024b), which provided detailed evaluation results for individual transfers, we restricted our comparison to the subset of apps from the dataset that are still functional. As discussed in Section 4, for TEMdroid (Zhang et al., 2024a), we could not obtain detailed experimental results even after contacting the authors. Due to the lack of available data, with respect to TEMdroid, we compared the average numbers for each category, which can be a threat to validity.
Our tool leverages Appium and the Appium UiAutomator2 Driver (Team, 2024) to interact with UI elements. This driver follows the WebDriver standard (Consortium, 2025), enabling a wide range of interactions with the app UI. However, since our tool builds on earlier approaches, we adopt a similarly limited yet well-curated set of interactions. While we support all actions covered by previous approaches—enhancing selector accuracy in the process—we do not provide full support for every possible action.
@@ 7. Related Works
The most relevant group of works targets transferring tests from one Android app to another, similar to LLMigrate. Behrang et al. (Behrang and Orso, 2018, 2019) and Lin et al. (Lin et al., 2019b) proposed ATM and CraftDroid, which rely on app analysis and NLP techniques to transfer tests across different Android apps within the same domain. AppFlow (Hu et al., 2018) is a machine learning-based approach that utilizes screen and widget classification to generate UI tests for an app using a library of existing tests. Liu et al. (Liu et al., 2022, 2024b) proposed adaptive semantic matching strategies for test transfer. Yu et al. (Zhang et al., 2024a) have recently proposed TEMdroid, a semantic matching-based approach for test transfer that leverages dynamic analysis and Siamese networks. Zhao et al. (Zhao et al., 2020) proposed FrUITeR, a framework for automatically evaluating the previous test transfer approaches. Mariani et al. (Mariani et al., 2021) presented a study on techniques for semantic matching of GUI events used by existing test reuse approaches. Khalili et al. proposed SemFinder (Khalili et al., 2024), an approach that assesses different configurations for mapping UI events across apps based on their textual information but does not focus on the test transfer problem as a whole. Mishra et al. (Mishra et al., 2023) extended SemFinder by incorporating visual information into the event mapping technique.
@@
Another group of recent works targets a slightly different problem, focusing on many-to-one UI test transfers. MigratePro (Zhang et al., 2024c) is a technique that aims to improve UI test transfer by generating a new test from multiple tests that have already been migrated to the target app from various source apps. MigratePro is not a transfer technique itself, and it improves tests transferred by an existing transfer technique. Future research can explore using tests transferred by LLMigrate as input for MigratePro to assess potential improvements. Another recent work, MACdroid (Zhang et al., 2024b), uses LLMs to create tests for a new target app using an abstract test logic created from multiple source tests targeting the same functionality on different apps. Note that compared to one-to-one transfer, many-to-one transfer is a less challenging problem due to the availability of more data, such as multiple different flows in various apps that can be more similar to the intended flow in the target app. However, in practice, there often are not multiple compatible tests for the same functionality available that can be used for transfer, which limits the practicality of these techniques. Furthermore, unlike LLMigrate, which employs multimodal LLMs and utilizes both visual and textual information for UI understanding tasks, MACdroid relies solely on textual data, which can lead to certain limitations, as discussed in Section 3. A more detailed comparison of the two approaches is not possible due to the unavailability of the implementation and artifacts relevant to this technique at the time of publication.
@@
Another group of related research focuses on bug reproduction in Android apps (Wang et al., 2024b; Zhao et al., 2019a, b). Note that although these works and test migration efforts both aim to execute a sequence of events in an Android app, the problems differ in two important aspects: (1) One of the biggest challenges of test migration comes from the differences in the flow of executed steps for a scenario between the source and target app. However, this issue does not exist in bug reproduction since the bug report belongs to the same app. (2) Test transfer also involves the challenge of accurately transferring and creating oracle events, which is a complexity that is not a part of the bug reproduction task.
Finally, several recent publications (Kang et al., 2023b, a; Feng and Chen, 2024; Huang et al., 2024; Liu et al., 2024a, c, 2023; Yu et al., 2023; Yoon et al., 2023; Wen et al., 2023; Ju et al., 2024) have explored the application of LLMs to advance mobile testing. However, none of these techniques address the automated transfer of existing usage-based tests across apps with similar functionality.
@@ 8. Conclusion and Future Work
This paper has presented LLMigrate, a technique that relies on multimodal LLMs for transferring usage-based UI tests across Android apps. In our extensive evaluation covering five app categories, LLMigrate was able to successfully transfer 97% of tests and reduce more than 90% of the total manual work required for writing UI tests.
Potential future areas of work can target expanding our technique, particularly for development across various platforms such as Web and iOS, which promises to yield significant time savings in end-to-end test development and maintenance. Another possible research direction is the study of how integrating the approaches (Chen et al., 2020) that enhance app accessibility can improve the applicability of techniques such as LLMigrate on non-accessible apps.
@@ 9. Data Availability
LLMigrate’s implementation and all of our research artifacts are available publicly (rep, 2025).
10. Acknowledgment
This work has been supported, in part, by award numbers 2106871, 2106306, and 2211790 from the U.S. National Science Foundation.